# Systems Security Foundations for Agentic Computing

### v1

@@ 3.4 Prompt Injection $\cong$ Dynamic Code Loading
then visit the GitLab page for that issue, see that the issue depends on another issue, and decide to summarize both to give the user a comprehensive answer. The distinction between (undesirable) prompt injection and (desirable) contextual task adjustment rests on the provenance of the input that triggered the change and the scope of the change.
@@
Prompt injection (benign or malicious) has a parallel in non-AI systems in the form of dynamic code loading. This is common in applications that support plugins or extensions (e.g., OS kernels, IDEs, web browsers) and has long posed significant challenges in terms of security. The typical solution is to introduce another security boundary between the main TCB and the dynamically loaded code and to use permissions or sandboxing to limit the access afforded to the new code. For example, web pages (probably the most popular platform for dynamic code loading, as each page load and subsequent interaction can potentially load additional Javascript code as web-page scripts) are secured through a number of sandboxes and associated security policies: Content Security Policy determines which third-party scripts can be loaded, Same-Origin Policy restricts the web page data accessible to a third-party script, <iframe> sandboxing isolates third-party content, and Subresource Integrity ensures that third-party scripts have not been modified from the time it was approved by the developer.
@@
Current agentic systems lack both components that provide security in web page dynamic code loading. First, information about the source (or provenance) of an instruction given to the agent is hard to assess (in contrast to the source of web-page scripts, most commonly loaded through explicit URLs). Second, sandboxing in the agent context is unavailable or probabilistic at best via mechanisms such as Instruction Hierarchy [wallace2024instructionhierarchytrainingllms]. An additional challenge to further drive the contrast with the web-page security model is that web pages are designed by a website owner, who is in charge of determining which code to include on the page and indirectly which code to dynamically load at some later time on the page. Agents are often given underspecified instructions and are expected to refine them with additional instructions discovered while working on the task at hand. This is a highly desirable functionality of agentic systems (not having to fully specify the task completely in the first prompt), but one for which security mechanisms are not readily available.
4 Case Studies of Real Attacks on Agentic Systems
@@
We present here 11 attacks against agentic systems to illustrate the variety of types of vulnerabilities that occur in practice. For each attack we also highlight the security principle(s) that were violated and the potential defenses.
@@
Microsoft Copilot Exfiltration. A vulnerability in Microsoft 365 Copilot allowed attackers to steal private user data, such as emails, by sending a malicious message containing a hidden prompt [wuest2024m365]. When a user interacted with this message (asking, for example, for a summary) via Copilot, a prompt-injection attack was triggered, letting the attacker take control of the agent. The compromised Copilot was then instructed to find sensitive information, encode it using “ASCII smuggling,” and embed it into a seemingly harmless hyperlink. When the user clicked this link, their data was secretly sent to the attacker.
@@
This attack violates the security principles of Least Privilege, Complete Mediation, and Secure Information Flow, as Copilot automatically performed unexpected actions sourced from a document of unknown origin (like searching for and exfiltrating data) without verifying each step with the user, failing to check that the AI’s operations were fully authorized. This vulnerability is similar to that of traditional code injection after a buffer overflow, where an adversarially crafted input allows the attacker to run code of their choice inside the victim process, coupled with insufficient access control.
@@
Devin AI Exposed Ports. The AI agent Devin comes with tool called expose_port, meant for testing, that was abused through an indirect prompt injection attack [wunderwuzzi2025devin]. An attacker hosted a malicious prompt on a website that, when visited by Devin, hijacked the agent. The compromised AI then started a local web server, exposing its entire file system, and used the expose_port tool to make this server publicly accessible online. The resulting URL was then sent to the attacker, granting them full access to Devin’s files.
This vulnerability violates the principles of Least Privilege (as the expose_port tool had excessive permissions, allowing it to expose any port, including one with access to the entire file system, rather than being restricted to only what was necessary for its intended function) and of Security Information Flow (as the agent accepted instructions received from an unknown origin). This vulnerability is similar to that of traditional code injection after a buffer overflow.
ChatGPT Long-Term Memory SpAIware. A vulnerability in the ChatGPT macOS app allowed for persistent data exfiltration by injecting malicious instructions into the app’s “Memories” feature [wunderwuzzi2024chatgpt]. This was accomplished through a prompt injection in an untrusted document or website, which caused the application to continuously send all of the user’s conversations to an attacker-controlled server. The data was exfiltrated by rendering an invisible image that included the user’s data as a parameter in the image URL.
This attack violates the principles of Trusted Computing Base (TCB) Tamper Resistance and Secure Information Flow because it stores data from an untrusted origin into the Memories storage (presumably trusted by the agent) and creates an unauthorized channel for sensitive information to be leaked from the application to an external, malicious server. In addition to the similarity with code injection after buffer overflow, this vulnerability also allows the attacker to achieve persistent control, which is one of the fundamental issues with dynamic code loading (Section 3.4).
Figure 2: OpenAI Operator Exploit Flow: Starts from an issue on a GitHub project and ends with PII being leaked to the attacker’s domain.
Amp AI Code Arbitrary Command Execution. Sourcegraph’s Amp AI coding agent allowed for arbitrary command execution by exploiting the agent’s ability to modify its own configuration file [wunderwuzzi2025amp]. Through a prompt injection attack, an adversary instructed the AI to alter its settings.json file, either by adding malicious commands to an allowlist for automatic execution or by adding an attacker-controlled server to the configuration, both of which led to running unauthorized code on the developer’s machine.
This attack fundamentally violates the principles of TCB Tamper Resistance (as the AI agent, a component of the trusted system, was able to modify its own security-critical configuration files, thereby compromising the integrity of the system’s security policies) and Secure Information Flow. In addition to the similarity with code injection after buffer overflow, this vulnerability has parallels with privilege escalation, where the attacker can modify the security configuration of their environment to gain higher access.
DeepSeek AI Account Takeover. A vulnerability in the DeepSeek AI platform led to a full account takeover by chaining a prompt injection with a Cross-Site Scripting (XSS) exploit [johann2024deepseek]. An attacker uploaded a malicious text file containing a base64 encoded JavaScript payload, which, when processed by a victim’s account, instructed the AI to decode and execute the script in the victim’s browser, stealing the userToken from localStorage and sending it to the attacker.
This attack violates the principles of Secure Information Flow (as it established an unauthorized covert channel to leak a sensitive session token from the user’s browser to an external, malicious endpoint, completely bypassing the platform’s intended data handling and security boundaries) and Complete Mediation (as the agent processed encoded data meant to bypass input filtering). Chained code injections are often used in traditional exploits to complete an attack.
Terminal DiLLMa. The “Terminal DiLLMa” attack hijacked a user’s terminal through LLM-powered command-line tools by using prompt injection to generate malicious ANSI escape sequences [wunderwuzzi2024terminal]. When a compromised tool, such as the proof-of-concept dillma.py, processed a malicious prompt, it output specially crafted ANSI codes that the terminal emulator executes, leading to unauthorized actions like clipboard manipulation or data exfiltration via DNS requests.
This attack highlights a violation of the Complete Mediation and Secure Information Flow principles. Insufficient or incorrect input validation and sanitization result in incompletely validated inputs which then trigger unwanted software behaviors, as it happened here.
ChatGPT Operator Prompt Injection. A prompt injection attack on the ChatGPT Operator led to the exfiltration of a user’s personally identifiable information (PII) by manipulating the agent through a malicious GitHub issue [johann2025chatgpt]. The attack began when the operator was prompted to investigate a GitHub issue containing a malicious combine tool, which, when clicked, redirected the agent to an attacker-controlled webpage. The compromised operator was then instructed to navigate to a settings page on another website (where the operator was already authenticated), to copy sensitive PII, and to paste it into a textbox on the attacker’s page, where it was immediately captured (see Fig. 2).
This exploit is a clear violation of Secure Information Flow, as the system fails to validate the origin of the request to perform various actions (navigating, clicking, copying, pasting) to ensure they are authorized by the user after the initial, legitimate prompt was given.
Devin AI Secret Leaks. The Devin AI agent was manipulated via indirect prompt injection to leak sensitive environment variables and secrets [wunderwuzzi2025devinLeak]. An attacker hosted a malicious prompt on a platform like GitHub, and when Devin was instructed to interact with it, the agent was tricked into using its native tools, such as the shell tool or the browsing tool, to exfiltrate data by sending it to an attacker-controlled server via commands like curl or by embedding it in a URL.
This vulnerability represents a failure of Secure Information Flow and Least Privilege principles, as it creates multiple unauthorized channels by unexpectedly executing tools (shell commands, browser navigation, markdown image rendering) for confidential data to be transmitted out of the agent’s secure environment.
Cursor AgentFlayer. A malicious Jira ticket was used to trick the AI-powered code editor, Cursor, into exfiltrating sensitive information [Simakov2025Jira]. The attack worked by creating a Jira ticket with a prompt that, while seemingly harmless, caused Cursor to leak repository secrets or even personal files, like AWS credentials, from the user’s local system. The author showed how simple changes in wording bypassed the AI’s built-in security measures. This exploit required, as a prerequisite, that a developer disabled the human-in-the-loop validation for the Jira MCP server or entirely enabled Cursor’s Auto Run mode (a form of YOLO mode for agents where confirmation prompts are minimized, often used in order to give the agents more freedom without annoying the developer).
This exploit violates the Secure Information Flow, Least Privilege, Complete Mediation, and TCB Tamper Resistance principles.
Claude Code Exfiltration. Claude Code allowed for data exfiltration via DNS requests [wunderwuzzi2025claude] (Fig. 3). The attack leveraged indirect prompt injection, where a malicious prompt hidden in a code file instructed Claude Code to read sensitive information, such as API keys from a .env file, and then used an allow-listed command like ping to send that data to an attacker-controlled server as part of a DNS query (via the inherent nslookup that occurs). The result was the leakage of sensitive information from the developer’s machine without their consent. Claude code had implemented human approval for many shell commands that could send datt to unknown domains, but mistakenly allowed ping to execute without human approval.
This violates secure information flow because the environment file contents leaked to an unknown domain. It also violated least privilege because the agent may not necessarily need uniform access to all shell commands with the ability to supply unrestricted arguments at all points in time.
Defense: The input arguments to tools such as ping and nslookup should be restricted to trusted or user-approved domains. Information-flow control for ML models (Section 5.3) and model architectures with built-in security controls (Section 5.5) could address such attack that chains multiple exploits.
AI ClickFix. Traditional social-engineering techniques were adapted to use against computer-use agents, in an attack called “AI ClickFix” [wunderwuzzi2025clickfix]. The attack involved tricking an AI agent into executing malicious code by presenting it with a series of instructions on a webpage. For instance, the agent was prompted to click a button which secretly copied a malicious command to the clipboard; then, the agent was instructed to open a terminal and paste the command from the clipboard into the terminal. The result was that the AI system was hijacked to execute arbitrary commands from untrusted web content.
This attack demonstrates that AI agents can be “socially engineered,” violating the principles of Human Weak Link, Least Privilege, and Secure Information Flow.
Defense: Introduce a hard security boundary between untrusted web content and privileged system tools. The agent should not be allowed to paste data from a browser clipboard directly into a system terminal without an user consent. This would restrict the agent’s functionality though and burdens the user with security decisions, effectively requiring the user to determine what is appropriate (visual) instruction and what is data (Section 5.1).
TABLE I: Examples of attacks and security principles they violate.

### v2

@@ 2.4 Dynamic Instruction Following
Prompt injection, more specifically indirect prompt injection, is often seen as a vulnerability of current AI-based systems. But more generally, prompt injection can be seen as a manifestation of a new class of functionality in AI, where the agent can adjust its task based on new instructions in its context. For example, using the GitLab task in Section 2.2, the agent may start with the aforementioned summarization request prompt, subsequently visit the GitLab page for that issue, identify that it depends on another issue, and decide to summarize both to give the user a comprehensive answer. The distinction between (undesirable) prompt injection and (desirable) contextual task adjustment rests on the provenance of the input that triggered the change and the scope of the change.
@@
From the security perspective, dynamic instruction following, which can be realized as prompt injection (benign or malicious), has a parallel in non-AI systems in the form of dynamic code loading. This is common among applications that support plugins or extensions (e.g., OS kernels, IDEs, web browsers) and has long posed significant challenges in terms of security. The typical solution is to introduce another security boundary between the main TCB and the dynamically loaded code and to use permissions or sandboxing to limit the access afforded to the new code. For example, web pages (probably the most popular platform for dynamic code loading, as each page load and subsequent interaction can potentially load additional Javascript code as web-page scripts) are secured through a number of sandboxes and associated security policies: Content Security Policy determines which third-party scripts can be loaded, Same-Origin Policy restricts the web page data accessible to a third-party script, <iframe> sandboxing isolates third-party content, and Subresource Integrity ensures that third-party scripts have not been modified from the time it was approved by the developer. Other systems, such as Android, actively discourage dynamic code loading [35].
@@
In the agentic space, dynamic instruction following is a key feature to steer task execution. For example, MCP descriptions are instructions brought in externally, as are Anthropic’s Claude Skills [6] and OpenClaw’s ClawHub add-ons [64]. Yet current agentic systems lack both of the components that provide security in web page dynamic code loading. First, information about the source (or provenance) of an instruction given to the agent is hard to assess (in contrast to the source of web-page scripts, most commonly loaded through explicit URLs over HTTPS). Second, sandboxing in the agent context is unavailable or probabilistic at best via mechanisms such as Instruction Hierarchy [96]. An additional challenge to further drive the contrast with the web-page security model is that web pages are designed by a website owner, who is in charge of determining which code to include on the page and indirectly which code to dynamically load at some later time on the page. Agents are often given underspecified instructions and are expected to refine them with additional instructions discovered while working on the task at hand. This is a highly desirable functionality of agentic systems (not having to fully specify the task completely in the first prompt), but one for which security mechanisms are not readily available.
3 Attacks on Agentic Systems: Case Studies
@@
We present here 11 attacks against agentic systems to illustrate the variety of types of vulnerabilities that occur in practice. For each attack we also highlight the security principle(s) that were violated and the potential defenses. For a refresher on the security principles we reference here, please see Appendix B. A summary of these attacks and corresponding violated security principles is shown in Table 1.
Microsoft Copilot Exfiltration.
@@
A vulnerability in Microsoft 365 Copilot allowed attackers to steal private user data, such as emails, by sending a malicious message containing a hidden prompt [73]. When a user interacted with this message (asking, for example, for a summary) via Copilot, a prompt-injection attack was triggered, letting the attacker take control of the agent. The compromised Copilot was then instructed to find sensitive information, encode it using “ASCII smuggling,” and embed it into a seemingly harmless hyperlink. When the user clicked this link, their data was secretly sent to the attacker.
@@
Violation: This attack violates the security principles of Least Privilege, Complete Mediation, and Secure Information Flow, as Copilot automatically performed unexpected actions sourced from a document of unknown origin (like searching for and exfiltrating data) without verifying each step with the user, failing to check that the AI’s operations were fully authorized. This vulnerability is similar to that of traditional code injection after a buffer overflow, where an adversarially crafted input allows the attacker to run code of their choice inside the victim process, coupled with insufficient access control.
@@
Table 1: Attacks of Section 3 and violated security principles.
@@
Devin AI Exposed Ports.
The AI agent Devin comes with tool called expose_port, meant for testing, that was abused through an indirect prompt injection attack [80]. An attacker hosted a malicious prompt on a website that, when visited by Devin, hijacked the agent. The compromised AI then started a local web server, exposing its entire file system, and used the expose_port tool to make this server publicly accessible online. The resulting URL was then sent to the attacker, granting them full access to Devin’s files.
Violation: This vulnerability violates the principles of Least Privilege (as the expose_port tool had excessive permissions, allowing it to expose any port, including one with access to the entire file system, rather than being restricted to only what was necessary for its intended function) and of Security Information Flow (as the agent accepted instructions received from an unknown origin). This vulnerability is similar to that of traditional code injection after a buffer overflow.
ChatGPT Long-Term Memory SpAIware.
A vulnerability in the ChatGPT macOS app allowed for persistent data exfiltration by injecting malicious instructions into the app’s “Memories” feature [74]. This was accomplished through a prompt injection in an untrusted document or website, which caused the application to continuously send all of the user’s conversations to an attacker-controlled server. The data was exfiltrated by rendering an invisible image that included the user’s data as a parameter in the image URL.
Violation: This attack violates the principles of Trusted Computing Base (TCB) Tamper Resistance and Secure Information Flow because it stores data from an untrusted origin into the Memories storage (presumably trusted by the agent) and creates an unauthorized channel for sensitive information to be leaked from the application to an external, malicious server. In addition to the similarity with code injection after buffer overflow, this vulnerability also allows the attacker to achieve persistent control, which is one of the fundamental issues with dynamic code loading (Section 2.4).
Figure 1: OpenAI Operator exploit flow: a prompt-injected GitHub issues page can steer the agent into an authenticated session (e.g., Gmail) and onward to an attacker-controlled site, resulting in personally identifiable information (PII) exfiltration.
Amp AI Code Arbitrary Command Execution.
Sourcegraph’s Amp AI coding agent allowed for arbitrary command execution by exploiting the agent’s ability to modify its own configuration file [77]. Through a prompt injection attack, an adversary instructed the AI to alter its settings.json file, either by adding malicious commands to an allowlist for automatic execution or by adding an attacker-controlled server to the configuration, both of which led to running unauthorized code on the developer’s machine.
Violation: This attack fundamentally violates the principles of TCB Tamper Resistance (as the AI agent, a component of the trusted system, was able to modify its own security-critical configuration files, thereby compromising the integrity of the system’s security policies) and Secure Information Flow. In addition to the similarity with code injection after buffer overflow, this vulnerability has parallels with privilege escalation, where the attacker can modify the security configuration of their environment to gain higher access.
DeepSeek AI Account Takeover.
A vulnerability in the DeepSeek AI platform led to a full account takeover by chaining a prompt injection with a Cross-Site Scripting (XSS) exploit [71]. An attacker uploaded a malicious text file containing a base64 encoded JavaScript payload, which, when processed by a victim’s account, instructed the AI to decode and execute the script in the victim’s browser, stealing the userToken from localStorage and sending it to the attacker.
Violation: This attack violates the principles of Secure Information Flow (as it established an unauthorized covert channel to leak a sensitive session token from the user’s browser to an external, malicious endpoint, completely bypassing the platform’s intended data handling and security boundaries) and Complete Mediation (as the agent processed encoded data meant to bypass input filtering). Chained code injections are often used in traditional exploits to complete an attack.
Terminal DiLLMa.
The “Terminal DiLLMa” attack hijacked a user’s terminal through LLM-powered command-line tools by using prompt injection to generate malicious ANSI escape sequences [75]. When a compromised tool, such as the proof-of-concept dillma.py, processed a malicious prompt, it output specially crafted ANSI codes that the terminal emulator executes, leading to unauthorized actions like clipboard manipulation or data exfiltration via DNS requests.
Violation: This attack highlights a violation of the Complete Mediation and Secure Information Flow principles. Insufficient or incorrect input validation and sanitization result in incompletely validated inputs which then trigger unwanted software behaviors, as occurred here.
ChatGPT Operator Prompt Injection.
A prompt injection attack on the ChatGPT Operator led to the exfiltration of a user’s personally identifiable information (PII) by manipulating the agent through a malicious GitHub issue [78]. The attack began when the operator was prompted to investigate a GitHub issue containing a malicious combine tool, which, when clicked, redirected the agent to an attacker-controlled webpage. The compromised operator was then instructed to navigate to a settings page on another website (where the operator was already authenticated), to copy sensitive PII, and to paste it into a textbox on the attacker’s page, where it was immediately captured (see Figure 1).
Violation: This exploit is a clear violation of Secure Information Flow, as the system fails to validate the origin of the request to perform various actions (navigating, clicking, copying, pasting) to ensure they are authorized by the user after the initial, legitimate prompt was given.
Devin AI Secret Leaks.
The Devin AI agent was manipulated via indirect prompt injection to leak sensitive environment variables and secrets [81]. An attacker hosted a malicious prompt on a platform like GitHub, and when Devin was instructed to interact with it, the agent was tricked into using its native tools, such as the shell tool or the browsing tool, to exfiltrate data by sending it to an attacker-controlled server via commands like curl or by embedding it in a URL.
Violation: This vulnerability represents a failure of Secure Information Flow and Least Privilege principles, as it creates multiple unauthorized channels by unexpectedly executing tools (shell commands, browser navigation, markdown image rendering) for confidential data to be transmitted out of the agent’s secure environment.
Cursor AgentFlayer.
A malicious Jira ticket was used to trick the AI-powered code editor, Cursor, into exfiltrating sensitive information [85]. The attack created a Jira ticket with a prompt that, while seemingly harmless, caused Cursor to leak repository secrets or even personal files, like Amazon Web Services (AWS) credentials, from the user’s local system. The author shows how simple changes in wording bypassed the AI’s built-in security measures. This exploit required, as a prerequisite, that a developer disabled the human-in-the-loop validation for the Jira MCP server or entirely enabled Cursor’s Auto Run mode (a form of YOLO mode for agents where confirmation prompts are minimized, often used in order to give the agents more freedom without involving the developer).
Violation: This exploit violates the Secure Information Flow, Least Privilege, Complete Mediation, and TCB Tamper Resistance principles.
Claude Code Exfiltration.
Claude Code enabled data exfiltration via DNS requests [79] (Figure 2 in Appendix A). The attack leveraged indirect prompt injection, where a malicious prompt hidden in a code file instructed Claude Code to read sensitive information, such as API keys from a .env file, and then used an allow-listed command like ping to send that data to an attacker-controlled server as part of a DNS query (via the inherent nslookup that occurs). The result was the leakage of sensitive information from the developer’s machine without consent. Claude code had implemented human approval for many shell commands that could send data to unknown domains, but mistakenly allowed ping to execute without human approval.
Violation: This violates secure information flow because the environment file contents leaked to an unknown domain. It also violated least privilege because the agent may not necessarily need uniform access to all shell commands with the ability to supply unrestricted arguments at all points in time.
Defense: The input arguments to tools such as ping and nslookup should be restricted to trusted or user-approved domains. Information-flow control for ML models (Section 5.3) and model architectures with built-in security controls (Section 5.5) could address such attacks that chain multiple exploits.
AI ClickFix.
Traditional social-engineering techniques were adapted to use against computer-use agents, in an attack called “AI ClickFix” [76]. The attack involved tricking an AI agent into executing malicious code by presenting it with a series of instructions on a webpage. For instance, the agent was prompted to click a button which secretly copied a malicious command to the clipboard; then, the agent was instructed to open a terminal and paste the command from the clipboard into the terminal. This resulted in the AI system being hijacked to execute arbitrary commands from untrusted web content.
Violation: This attack demonstrates that AI agents can be “socially engineered,” violating the principles of Human Weak Link, Least Privilege, and Secure Information Flow.
Defense: Introduce a hard security boundary between untrusted web content and privileged system tools. The agent should not be allowed to copy data from a browser clipboard directly into a system terminal without user consent. This would, however, restrict the agent’s functionality and burdens the user with security decisions, effectively requiring the user to determine what is an appropriate (visual) instruction and what is data (Section 5.1).
Table 2: Comparison of agentic AI safety systems discussed in Section 4.
Work Challenges Addressed
Details
TCB Assumption
How probabilistic components are handled, e.g., fine-tuning, external deterministic monitors, isolation, crypto/ACL primitives, or attention-based analysis.
Policy Support
Source or derivation mechanism for task-time permissions and policies: generated by an LLM, automated via non-LLM heuristics or external policy inputs, user-mediated via interactive approvals, user-defined via explicit configuration, or N/A when policies are static/pre-defined.
Trust Boundary
Where the security boundary is placed: Model, Detector, Tool level, IR/Plan level, App/System level, Inter-agent level, Action level, HTTP request layer, or Decision level.
Instruction Following
Mechanisms addressing prompt injection as dynamic instructions to the agent.
Probabilistic TCB Dynamic Policy Fuzzy Trust Boundary Dyn. Instr. Following
f-secure [101] Security monitor, planner/executor separation Generated (task) System level (IFC) Monitor filters untrusted input
FIDES [23] Deterministic monitor, security labels User-defined (IFC labels) Tool level, IR Data hiding, IFC labels
IsolateGPT [103] Hub-and-spoke isolation, hub as kernel User-mediated (cross-app) App level (spoke isolation) Execution isolation
ACE [48] Trusted abstract plan, static analysis N/A (abstract/concrete decoupling) Plan level (IR) Static IFC on structured plan
CaMeL [25] Control flow isolation Generated (trusted query) IR Plan from trusted data
Progent [84] Deterministic monitor
User-defined (JSON) Tool level Static policy
Progent-LLM [84] Prompting
Generated (full context) Tool level Prompting
SAGA [90] Central Provider, crypto tokens
User-defined (access control) Inter-agent level Access control tokens
ShieldAgent [19] Predicate classification, action-based safety policy model (ASPM) Automated (policy documents) Tool level (action-based) Policy verification via ASPM
GuardAgent [105] Guard agent (LLM + code execution) Generated (task) Action level Code-based guardrails
LlamaGuard [43] Fine-tuning for safety classification N/A Detector N/A (content moderation)
SecAlign / IH [18, 96] Fine-tuning (adversarial DPO)
N/A Model Adversarial training
StruQ [16] Fine-tuning for instruction format N/A Model Instruction format enforcement
DataSentinel [52] Fine-tuning for instruction detection N/A Detector Known-answer detection
ceLLMate [54] Browser extension, HTTP mediation Automated + website policies HTTP request layer Sandboxing blast radius
SkillFence [41] Browser extension, phonetic analysis N/A Skill invocation level N/A (voice confusion)
MindGuard [98] Attention-based DDG (non-invasive) N/A (policy-agnostic) Decision level (DDG) Detection and attribution
RTBAS [113] LM-as-judge + attention saliency Generated (task) Tool level (IFC) Dependency screeners
DRIFT [49] Secure Planner + Dynamic Validator Generated (task) Control + data level Injection Isolator
NeMo Guardrails [61] Colang runtime, rail engine User-defined (Colang) Input/output/dialog rails Programmable rails
Jamshidi et al. [45] RSA signing, LLM semantic vetting N/A Tool descriptor level Manifest signing + vetting
VeriGuard [56] Formal verification, runtime monitor Generated (task) Action level Verified policy + runtime check
4 Current Approaches
We provide an overview of existing efforts to secure agentic systems, with a concise summary in Table 2.
4.1 General AI Security
In the past decade, there has been ample research on formally and rigorously verifying the correctness of AI-based systems in general, and neural networks in particular [1]. The various techniques define a property of interest, e.g., is the AI classifier $\epsilon$-robust to noise, and translates this check to a mathematical query that is then solved by off-the-shelf optimizers. On one hand, methods that are both sound (always correct when they claim safety) and complete (always find a violation, if it exists and resources permit) check if the required property holds for every possible input within a domain of interest. Such methods include MILP [94], SAT [59], SMT [46, 47], and are typically geared towards AI models with piecewise-linear activation functions [14, 13, 46]. In some cases, these also include branch-and-bound (B&B) methods, e.g., [109, 97, 13]. Alternatively, some approaches sacrifice completeness (and hence, may not find existing counterexamples) but gain performance. Such techniques include IBP [57, 38], Zonotopes [57, 32], DeepZ [86], DeepPoly [87, 88]), convex relaxations [99], and SDP-based verification [69, 68]), as well as certificate-based techniques (Lyapunov/barrier functions [15, 67]).
Finally, complementing methods are lightweight verification during runtime such as [2].
Robustness can also be assessed with gradient attacks [33, 53] and testing techniques [66, 11].
4.2 Agentic AI Security
Compared to standalone models, verifying the security of agentic AI systems is substantially more challenging for multiple reasons. First, agents routinely invoke external, third-party tools whose outputs are not under the developer’s control and may be adversarial. Second, interaction is typically iterative: intermediate tool outputs are incorporated back into the agent’s context and can influence downstream planning and actions, creating long-range dependencies and compounding risk over time. Third, both the user interface and the tool interface are often mediated through natural language (tool descriptions, user requests, and free-form tool results), making the “API surface” less structured and more susceptible to instruction smuggling and indirect prompt injection attacks [39, 51].
Threat Landscape.
Recent work has systematically characterized the attack of LLM agents. Attacks can be broadly categorized as agent-based, which manipulate internal components such as instructions [112], memory or knowledge bases [20, 116, 104, 22], as well as tool libraries [31, 107]. On the other hand, environment-based attacks exploit vulnerabilities in the external environment the agent interacts with, such as injecting malicious HTML elements [106] or deceptive pop-ups [111]. These attacks can lead to severe consequences, including privacy breaches [50], financial losses [5], and life-threatening failures [20].
Dynamic Monitor-Based Defenses.
One line of work preserves the expressive "plan-act-observe-replan" loop but requires continuous enforcement because tool results can modify the effective context seen by the planner. f-secure uses a context-aware pipeline with structured executable plans and an explicit security monitor that filters untrusted inputs entering the planning process [101]. Similarly, FIDES applies dynamic information-flow control to agent planning by tracking integrity/confidentiality labels and enforcing policies at each step [23]. IsolateGPT introduces an architecture that isolates execution environments across applications, requiring user intervention for potentially dangerously cross-app communications [103]. In these architectures, sandboxing/isolation and runtime monitoring are naturally complementary: isolation limits the potential damage of using malicious tools, while monitoring constrains how untrusted data can steer subsequent decisions. Frameworks such as CAPSEM [36] and Invariant Guardrails [9] provide the enforcement mechanisms for applying a variety of security policies, from PII leakage detection to tool-call guardrails to dataflow controls, configurable via expressive policy languages [37].
Planning-First (Static) Approaches.
An alternative paradigm intentionally reduces feedback from untrusted tool outputs into future control flow, enabling more principled ahead-of-time security reasoning. ACE [48] separates query processing into abstract plan-generation from trusted user input, concrete plan instantiation, and isolated execution, then statically checks information flow constraints over the resulting structured plan. CaMeL [25] extracts control and data flows from trusted user queries and employs a custom interpreter to prevent untrusted data from affecting program flow. These approaches sacrifice some interactivity to minimize the need for monitoring the tool-response channel and strengthen enforceable security boundaries.
Privilege Control and Policy Languages.
Drawing inspiration from classical access control systems [28, 24] and cloud identity and access management (IAM) policies [3, 55, 34], recent work explores programmable privilege control for agents. Progent [84] introduces a domain-specific language for expressing fine-grained, tool-call-level policies that specify which actions are permissible, under what conditions, and with what fallback behaviors. Policies can be manually specified for deterministic security guarantees or automatically generated and updated by LLMs to adapt to task-specific requirements. This approach enforces the principle of least privilege: blocking unnecessary and potentially malicious tool calls while permitting those essential for task completion. Evaluation on benchmarks including AgentDojo [26] and ASB [108] shows substantial reductions in attack success rate (e.g., from $41.2\%$ to $2.2\%$) while maintaining utility.
Guardrail Agents and Policy Reasoning.
While traditional guardrails focus on content moderation for LLMs as models (e.g., LlamaGuard [43], LlavaGuard [40]), they fail to address the complexities of action sequences where vulnerabilities emerge over time [105]. ShieldAgent [19] addresses this gap by introducing a guardrail agent that shields other agents via verifiable safety policy reasoning. It constructs an action-based safety policy model (ASPM) by extracting verifiable rules from policy documents (e.g., EU AI Act, corporate handbooks) and organizing them into probabilistic rule circuits. During inference, ShieldAgent retrieves relevant circuits for the invoked action, generates verification plans using specialized tools, and performs formal verification via model checking before making probabilistic guardrail decisions. This approach achieves high accuracy while reducing computational overhead compared to naive rule traversal.
Model-Level Defenses.
Complementing system-level and user-level approaches, model-level defenses aim to make LLMs inherently more robust to prompt injection. These include fine-tuning approaches that train models to ignore injected prompts [16, 18, 96]. Detection-based defenses such as DataSentinel [52] use a secondary LLM to identify contaminated inputs via known-answer detection (KAD), achieving a near-perfect accuracy against existing attacks. However, KAD schemes contain a structural vulnerability [21]: since the detection instruction and secret key share the same context window as the potentially malicious input, adaptive attacks can extract and return the secret key while still executing the injected task, fundamentally undermining the defense. These defenses operate at a different level than system-level privilege control and can work in synergy model defenses protect the core reasoning while system defenses safeguard the execution boundary between the agent and the external tools.
At a high level, these approaches represent recurring trade-offs: dynamic agent loops offer adaptivity but demand persistent monitoring of tool responses and context evolution; planning-first architectures sacrifice interactivity to strengthen enforceable security boundaries; privilege control policies require careful specification but enable deterministic guarantees; and guardrail agents add verification overhead but provide explicit policy compliance with interpretable explanations.